Walter and his parents are a little unusual. If he does an acceptable job of
doing his chores ,he gets paid $3.33 fo that day. If he does an outstanding
job, he gets paid $3 more. During one 10 day period, Walter received $42.30 for
his work. How many days did he do an acceptable job?  How many days did he do
an exceptional job?

Answer by Edwin McCravy(20086)


Let the number of days he does an acceptable job = x.
then the number of days he does an outstanding job = 10-x

He got $3.33 for each of the x days, which amounted to $3.33x
He got $3 more for each of the 10-x days, or $3.33+$3 or $6.33,
which amounted to $6.33(10-x)


Get the equation from this:

(The money he got for all x days when he did an acceptable job) 

                       PLUS

(The money he got for all 10-x days when he did an outstanding job)

                      EQUALS

                      $42.30

So the equation is

               $3.33x + $6.33(10 - x) = $42.30
                 3.33x + 6.33(10 - x) =  42.30
Drop the zero
on 42.30, making 
it 42.3
                 3.33x + 6.33(10 - x) =  42.3
Distribute
                 3.33x + 63.3 - 6.33x =  42.3
combine x terms                  
                        -3.00x + 63.3 =  42.3
Drop the zeros
on -3.00x, making
it -3x

                           -3x + 63.3 =  42.3 
subtract 63.3 
from both sides:
                                -63.3 = -63.3
                       ----------------------
                           -3x        = -21

Divide both sides by -3

                           -3x          -21
                          -----       = ----
                           -3            -3

Cancel the -3's on the
left and divide on the
right
                             x        =   7            

The number of days he did an acceptable job = x = 7 days
The number of days he did an outstanding job = 10-x = 10-7 = 3 days
 
Edwin